OpenStudy (pandora):

Find the limit as x goes to 0 of 1-cosx/sin x

Parth (parthkohli):

Use L'Hospital's Rule.

myininaya (myininaya):

Or you can use the following instead: Recall the following: \[\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=1 ; \lim_{x \rightarrow 0}\frac{\cos(x)-1}{x}=0\]

Parth (parthkohli):

\[\lim_{x \to 0}{-\sin (x) \over \cos(x)} \]After differentiating the numerator and denominator, we get the above.

OpenStudy (pandora):

But how do you use the second part of that when sin is in the denominator, not just x

Parth (parthkohli):

We may now plug-in 0 for \(x\).

myininaya (myininaya):

Derivative of (-cos(x)) is .....?

Parth (parthkohli):

Oh, my.

Parth (parthkohli):

I am messing up questions: one-by-one.

Parth (parthkohli):

\[\lim_{x \to 0} {\sin(x) \over \cos(x)} \]

Parth (parthkohli):

Sorry again @Pandora @myininaya

OpenStudy (pandora):

No problem but I'm still lost sorry is there any other way to explain this

Parth (parthkohli):

Graphing can help, yes.

Parth (parthkohli):

Note: never ever try plugging and checking values if it's not a polynomial.

myininaya (myininaya):

Well you can use the properties I mentioned: \[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{\sin(x)} =\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x} \cdot \frac{x}{\sin(x)}\] Recall x/x=1 I just multiplied your expression by 1 (a funky one but it is still a one nonetheless)

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x} \cdot \lim_{x \rightarrow 0}\frac{x}{\sin(x)}\]

myininaya (myininaya):

You should be able to evaluate both of those limits given the properties I gave you.

Parth (parthkohli):

If you still won't get that, then check out the grap here: http://www.wolframalpha.com/input/?i=%281-cosx%29%2Fsin+x

Parth (parthkohli):

graph*

OpenStudy (pandora):

Oh wow thanks so much you guys. The lightbulb finale went on. Thanks for the help

myininaya (myininaya):

Np.